OpenStudy (anonymous):

Determine whether the following has a maximum or minimum value and then compute that value: y = 2x2 + 10x - 4

OpenStudy (anonymous):

is that 2 times 2 or no

OpenStudy (anonymous):

do you have any answer choices

OpenStudy (anonymous):

yeah but i found it i need help with this one Determine the quadratic equation for the parabola: range: {y:y ≤ 9}; x-intercepts: -2 and 4

OpenStudy (anonymous):

have you already tackled derivatives?

OpenStudy (anonymous):

no im taking online classes so some of the stuff i dont know

OpenStudy (anonymous):

lol same here wait a few ill get you an answer

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

thankyouuuu

OpenStudy (anonymous):

ok here's a review. to find the maximum and minimums of the value of a function, what we commonly do is we first find the critical points

OpenStudy (anonymous):

now what are critical points? these are points where the derivative of the function of f'(x)=0

OpenStudy (anonymous):

so suppose if we have any constant c and I plug this in to f'(x) and get 0 or if f'(c)=0 then the function will have a critical point at x=c

OpenStudy (anonymous):

wow im stumped lol heressometimes try to put you question in google or yahoo to get a

OpenStudy (anonymous):

fasster answer

OpenStudy (anonymous):

now this critical point could either be a minimum point or a maximum point

OpenStudy (anonymous):

k cuase this seems to complicated

OpenStudy (anonymous):

ok ok :)) I'll make it simplier then

OpenStudy (anonymous):

lol when everyone confuses me i just look on google and yahoo for the answer lol

OpenStudy (anonymous):

first find f'(x)

OpenStudy (anonymous):

here we have f'(x)=4x+10

OpenStudy (anonymous):

yeah they got this site called expert.com and it has alot of answers you just type in the proplem

OpenStudy (anonymous):

haha that even confuses me hey anony try my question lol

OpenStudy (anonymous):

If you know the x intercepts of the "equation" you can reason towards the form asked for.

OpenStudy (anonymous):

ok ok :))

OpenStudy (anonymous):

For example, you know that x=-2 will make f(x)=0, that is f(-2)=0

OpenStudy (anonymous):

btw it has a minimum which is at point x=-5/2

OpenStudy (anonymous):

Moreover, you know that x=4 will make the function attain a 0value. f(4)=0

OpenStudy (anonymous):

This leads to the equation: (x-4)*(x+2), the 2 factors are the "roots" of the equation - where the value of the function is 0.

OpenStudy (anonymous):

Which results in x^2+2x-4x-8 --> f(x)=y=x^2-2x-8